Jim Shorts Company makes clothing for schools. Sales in 20X1 were $4,880,000. Assets were as follows Cash Accounts receivable Inventory $ 116,000 822,000 427,000 Net plant and equipment517,000 $ 1,882,000 Total assets a. Compute the following: (Round your answers to 2 decimal places.) Turnover Ratios Accounts receivable 1. turnover 2. Inventory turnover 3.Fixed asset turnover 4. Total asset turnover times times times timesb. In 20X2, sales increased to $5,470,000 and the assets for that year were as follows: Cash Accounts receivable Inventory Net plant and equipment $ 116,000 967,000 990,000 517,000 $ 2,590,000 Total assets Compute the following: (Round your answers to 2 decimal places.) Turnover Ratios 1. Accounts receivable turnover 2. Inventory turnover 3. Fixed asset turnover 4. Total asset turnover times times times times
Answer #1

a.compute the following:

turnover ratios
1. accounts receivable turnover 5.94 times
2. inventory turnover 11.43 times
3. fixed asset turnover 9.44 times
4. total asset turnover 2.59 times

working:

accounts receivable turnover = sales / accounts receivable

=>4,880,000/822,000

=>5.94

inventory turnover =>sales / inventory

=>4,880,000 / 427,000

=>11.43

fixed asset turnover = sales / net plant and equipment

=>4,880,000 / 517,000

=>9.44 times.

total asset turnover

=>sales / total assets

=>4,880,000 / 1,882,000

=>2.59 times.

b.compute the following:

turnover ratios
1. accounts receivable turnover 5.66 times
2. inventory turnover 5.53 times
3. fixed asset turnover 10.58 times
4. total asset turnover 2.11 times

working:

accounts receivable turnover =5,470,000 / 967,000

=>5.66 times

inventory turnover = 5,470,000 / 990,000

=>5.53

fixed asset turnover = 5,470,000/517,000

=>10.58 times

total asset turnover =5,470,000 / 2,590,000

=>2.11 times.